roots of quation log2(n2-4n+5)=n-2 are

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

log_{2}(n^{2}-4n+5) = n-2
f(n) = n^{2}-4n+5
is always +ve for all values of n. So domain is OK.
Now, if we draw the graph
f_{1}(n) = log_{2}(n^{2}-4n+5)
For n = 2, it touches the x – axis.
And for
f_{2}(n) = n-2
For n = 2, it touche the x – axis.
From the grapg it is clear that n = 2 is the solution.